OpenStudy (anonymous):

Find the linearization of f(x)=(1+x)^k at x=0. Your answer will be in terms of x and k. Could you take a look at this one too when ya get the chance also @misty1212? Specifically when I'm deriving (1+x)^k, would it equal k(1+x)^(k-1)?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

course we didn't do so well with the last one, even though i am sure it was right

OpenStudy (anonymous):

No you were right it was asking for the formula, not the answer :(

OpenStudy (anonymous):

I didn't realize hahah

OpenStudy (anonymous):

Thanks for the help I really do appreciate it. I suck at this stuff.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

I have 1^k+k(1+x)^(k-1) But what would the (x-a) part equal of the equation? x=0, but what is a?

OpenStudy (misty1212):

x is the variable, a is the zero

OpenStudy (misty1212):

the point is \((0,1)\)

OpenStudy (misty1212):

you get \[f'(x)=k(1+x)^{k-1}\] which at \(x=0\) gives the slope of \\(k\)

OpenStudy (misty1212):

and the line of \[y-1=k(x-0)\] or \[y=kx+1\]

OpenStudy (anonymous):

Got it! Thanks again

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]